# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ G.1 Misspecified linear bandits
$\leq 2\|M\|_{\max}\delta T+\sum_{t=T_{1}+1}^{T}\mathbb{E}\left[\left(r_{t,a^{\star}_{t}}-r_{t,a^{\pi}_{t}}\right)\mathds{1}_{\{\mathcal{E}\}}\right]$
Recall that $\|M\|_{\max}=\Theta(1)$ and define $\mathcal{L}_{mnT}=\log(1+mnT)$. Then, we select $\delta=1/T$, and after substitution of $\epsilon_{\max}$ from (47) and $\delta$ in (46), we have:
$R_{2}^{\pi}\leq C_{3}\Big{(}1+\sqrt{r(m+n)}\Big{(}\sqrt{T\mathcal{L}_{mnT}}+\frac{\mu^{6}\kappa^{4}r^{3}(m+n)T}{\omega_{\min}(m\wedge n)\sqrt{mn}T_{1}}\mathcal{L}_{mnT}^{3}+\sqrt{r(m+n)}\Big{)}\mathcal{L}_{mnT}^{2}\Big{)}$ (48)
provided that $T_{1}\geq\frac{c\mu^{4}\kappa^{2}r^{2}(m+n)}{\omega_{\min}mn}\log^{3}((m+n)T^{2})$ for some universal constant $C_{3},c>0$ and $T_{1}\leq T$.
Step 3: (Optimizing $T_{1}$). Observe that the upper bound $R^{\pi}_{1}$ scales as $T_{1}$ (see (45)) while the third term in the upper bound of $R^{\pi}_{2}$ scales as $1/T_{1}$ (see (48)). Therefore, to obtain a tight regret bound on $R^{\pi}(T)$, we need to balance these two terms that depend on $T_{1}$ with a proper choice of $T_{1}$. We can easily verify that a good choice is:
$T_{1}=\sqrt{\frac{C_{3}}{C_{1}}}\mu^{3}\kappa^{2}r^{7/4}\frac{(m+n)^{3/4}}{\sqrt{\omega_{\min}}(m\wedge n)^{1/2}(mn)^{1/4}}\sqrt{T}\log^{5/2}\left(1+mnT\right).$ (49)
Note that the condition on $T_{1}$ is valid for $T=\widetilde{\Omega}\left(\mu^{2}r^{1/2}\frac{(m+n)^{3/2}}{\omega_{\min}(mn)^{3/2}},\mu^{6}\kappa^{4}r^{7/2}\frac{(m+n)^{3/2}}{\omega_{\min}(m\wedge n)\sqrt{mn}}\right)$. Substituting this $T_{1}$ back into Equation (48) and adding $R_{1}^{\pi}$ gives the regret upper bound given in the statement of Theorem 8.1 after noting that $\omega_{\min}=m\min_{i}\rho_{i}/(mn)$. Finally, note that for $T$ smaller than the required value, the trivial regret upper bound of $T\|M\|_{\max}$ is smaller than the regret in Theorem 8.1.
Appendix H Reduction to an almost low-dimensional linear bandits and its limitations
@@
and note that $M_{i,j}=\langle\phi_{i,j}^{\textup{ext}},\theta^{\textup{ext}}\rangle$ according to (3). Moreover, one can readily show that, for any $(i,j)\in[m]\times[n]$.
@@
$\|\theta_{\ell}\|_{2}\lesssim\frac{\mu^{4}\kappa^{2}r^{3}(m+n)}{T_{1}\omega_{\min}\sqrt{mn}}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right)$
@@
From this observation, together with the reformulation $M_{i,j}=\langle\phi_{i,j}^{\textup{ext}},\theta^{\textup{ext}}\rangle$, one can claim a reduction from a low-rank bandit model to an almost low-dimensional linear bandit (Valko et al., 2014; Jun et al., 2019; Kocák et al., 2020; Kang et al., 2022). More precisely, the sub-vector corresponding to the last $(m-r)(n-r)$ components of $\theta$ (i.e., those corresponding to $\theta_{\ell}$) do not have a significant impact on the rewards, hence reducing the effective dimension of this linear bandit to $r(m+n)-r^{2}$.
Similarly to Algorithm 2 and the regret analysis of Theorem 8.1, we can show:
@@
$\Lambda=\mathrm{diag}(\overbrace{\lambda,\dots,\lambda}^{d},\overbrace{\lambda_{\perp},\dots,\lambda_{\perp}}^{mn-d}),\qquad\beta=\sqrt{2\log\left(\frac{10Tmn}{\delta}\log(1+T)\right)}+B_{2}\sqrt{\lambda}+B_{\ell}\sqrt{\lambda_{\perp}}.$
Then, Algorithm 2 with SupLinUCB parameters $T-T_{1},\beta(\delta),\Lambda$ has an expected regret that satisfies $R^{\pi}_{T}=\widetilde{O}(\mu^{2}\kappa r^{3/2}\sqrt{\frac{m+n}{\omega_{\min}\sqrt{mn}}}\sqrt{T})$.
The proposition follows from a similar analysis as in the proof of Theorem 8.1. Instead of Theorem G.1, we combine the analysis of SupLinUCB [^15] Here we used SupLinUCB as given in (Chu et al., 2011) for the analysis. from (Chu et al., 2011), Lemma 38 in (Kocák et al., 2020), Lemma 3 (Jun et al., 2019), and Lemma 11 in (Abbasi-Yadkori et al., 2011) to show that the second phase of this algorithm has regret as follows:
${\cal R}^{\pi}(T)\lesssim\sqrt{dT}\left(\sqrt{\log\left(eTK/\delta\right)}+B_{2}\sqrt{\lambda}+B_{\ell}\sqrt{\lambda_{\perp}}\right)\sqrt{\log\left(1+T/\lambda\right)\log T}$
@@
(i) First, the above reduction does not take advantage of the delocalization of the subspace recovery error, i.e., of the fact that this error spreads out along $m+n$ dimensions. This delocalization property is quantified through our subspace recovery error guarantees in the two-to-infinity norm, and it has subtle but important implications. Indeed, it can be leveraged to perform a reduction to an almost low-dimensional linear bandit that is also almost sparse, in the sense that $\|\theta_{\ell}\|_{\infty}\lesssim\|\theta_{\ell}\|_{2}/\sqrt{mn}$. A reduction to a misspecified linear bandit leverages this structure and allows us to devise algorithms with tighter regret guarantees.
@@
(iii) We finally note that a reduction to misspecified linear bandit is more appealing computationally because the resulting feature vectors are of dimension $r(m+n)-r^{2}$. In the case of the reduction to an almost low-dimensional linear bandit, these vectors are of dimension $mn$. This means that instead of having to invert [^16] Indeed, almost all existing algorithms for linear and contextual linear bandits require inverting the features matrix during a regression step that involves least squares estimation (Lattimore & Szepesvári, 2020). $mn\times mn$ matrices, we will only have to invert $(r(m+n)-r^{2})\times(r(m+n)-r^{2})$ matrices.
@@ Appendix I Additional discussions
I.1 Low-rank matrix bandits
Our framework can be applied to adress a related and simpler problem known as low-rank matrix bandits. In this scenario, the context set remains constant in every round $\mathcal{X}_{t}=\mathcal{X}=\{x_{1},\dots,x_{K}\}$, and the learner must choose two arms, $i_{t}\in[m]$ and $j_{t}\in[n]$, at each time step $t\in T$. Unlike the contextual low-rank bandits setting, where $i_{t}$ is sampled according to a distribution $\rho$, in low-rank matrix bandits, the learner can decide which row $i_{t}$ is sampled next.
@@ I.1 Low-rank matrix bandits
Similarly to Theorem 8.1, it can be proven that such an algorithm would achieve a regret of at most $\widetilde{O}(\mu^{3}\kappa^{2}r^{7/4}(m+n)^{3/4}(mn)^{1/4}\sqrt{T/(m\wedge n)})$ under an appropriate condition on $T$ (see Theorem 8.1). When $M$ is homogeneous, the regret scales as $\widetilde{O}((m+n)^{3/4}\sqrt{T})$.
Most existing work on low-rank bandits lacks minimax regret guarantees. (Katariya et al., 2017) introduced Rank1Elim, achieving regret $\widetilde{O}((m+n)\log(T)/\tilde{\Delta}_{\min})$ [^17] Here $\tilde{\Delta}_{\min}$ corresponds to a notion of minimum gap different from the standard one - see (Katariya et al., 2017). for a rank-1 reward matrix. (Kveton et al., 2017) extended this result to rank $r$, with additional strong assumptions on the reward matrix in their LowRankElim algorithm. (Trinh et al., 2020) provided the first algorithm with asymptotically optimal instance-dependent regret for rank-$1$ bandits. (Bayati et al., 2022) proposed an algorithm enjoying a minimax regret guarantee. However, their upper bounds scale at least with $mn$ through the constant $C_{2}$ in the worst case (see Theorem 2 in (Bayati et al., 2022)), and requires tuning a filtering resolution $h$, making their guarantees hard to evaluate. (Stojanovic et al., 2023) presented SME-AE algorithm, leveraging the entry-wise matrix estimation guarantees to obtain a regret upper bound of order $\widetilde{O}((m+n)(\bar{\Delta}/\Delta_{\min}^{2})\log^{3}(T))$ where $\bar{\Delta}$ is the average gap, and $\Delta_{\min}$ is the minimum gap.
While these gap-dependent bounds require strong assumptions, e.g., (Stojanovic et al., 2023) assume that $\Delta_{\max}/\Delta_{\min}\leq\zeta$ where $\Delta_{\max}$ is the maximum gap, we provide the first algorithm with a minimax regret of order $\widetilde{O}((m+n)^{3/4}\sqrt{T})$ without such assumptions.
I.2 Improved max-norm guarantees
@@
$\|M-\widehat{M}\|_{\max}\lesssim\mu^{11/2}\kappa^{2}r^{3/2}\|M\|_{\max}\frac{(m+n)^{3/2}}{\sqrt{T\omega_{\min}mn}\left(m\wedge n\right)}\log^{3/2}\left(\frac{(m+n)T}{\delta}\right)$
with probability larger than $1-\delta$ when $T\gtrsim\frac{\mu^{4}\kappa^{2}r^{2}\left(m+n\right)}{\omega_{\min}mn}\log^{3}\left((m+n)T/\delta\right)$.
@@
with probability larger than $1-\delta$ under the assumptions of Proposition D.2. When $T$ is large enough to ensure that the first term in the error bound is larger than the second, the scaling in $\mu$ and $r$ of the max-norm error bound is improved from $\mu^{11/2}r^{3/2}$ to $\mu r^{1/2}$. We also note that the dependency in $\kappa$ and $\|M\|_{\max}$ is removed. This may appear suprising, but note that the term $O\left(\frac{\log^{3}T}{T}\right)$ still depends on $\kappa$ and $\|M\|_{\max}$ : more samples are required to achieve the stated scaling when $\kappa$ and $\|M\|_{\max}$ are large.

### v2

@@ G.1 Misspecified linear bandits
$\leq 2\|M\|_{\max}\delta T+\sum_{t=T_{1}+1}^{T}\mathbb{E}\left[\left(r_{t,a^{\star}_{t}}-r_{t,a^{\pi}_{t}}\right)\mathds{1}_{\{\mathcal{E}\}}\right].$
Define $\mathcal{L}_{mnT}=\log(1+mnT)$. Substituting $\epsilon_{\max}$ with its upper bound in (49) and selecting $\delta=1/T$ in (48) gives:
$R_{2}^{\pi}\leq C_{3}\sqrt{r(m+n)}\Big{(}\sigma\sqrt{T\mathcal{L}_{mnT}}+T\frac{L^{2}\kappa^{3}\mu^{2}r(m+n)}{\sigma_{r}T_{1}\omega_{\min}(m\wedge n)}\mathcal{L}_{mnT}^{3}+\|M\|_{\max}\sqrt{r(m+n)}\Big{)}\log^{2}\left(T(m+n)\frac{L^{2}}{\sigma^{2}}\right)$ (50)
provided that $T_{1}\geq c\frac{L^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left((m+n)T^{2}\right)$, for some universal constant $C_{3},c>0$ and $T_{1}\leq T$.
Step 3: (Optimizing $T_{1}$). Observe that the upper bound of $R^{\pi}_{1}$ scales as $T_{1}$ (see (47)) while the second term in the upper bound of $R^{\pi}_{2}$ scales as $1/T_{1}$ (see (50)). Therefore, to obtain a tight regret bound on $R^{\pi}(T)$, we need to balance these two terms that depend on $T_{1}$ with a proper choice of $T_{1}$. We can easily verify that an optimal choice is:
$T_{1}^{\star}=\sqrt{\frac{C_{3}}{2}}\frac{\mu\kappa^{3/2}L}{\sqrt{\|M\|_{\max}}}\frac{r^{3/4}(m+n)^{3/4}}{\sqrt{\sigma_{r}\omega_{\min}(m\wedge n)}}\sqrt{T}\log^{3/2}\left(1+mnT\right)\log\left(T(m+n)\frac{L^{2}}{\sigma^{2}}\right).$
Next, note that we can use that $\|M\|_{\max}/\sigma_{r}(M)\leq\mu^{2}\kappa r/\sqrt{mn}$ from Lemma B.8 and the fact that $\omega_{\min}=\rho_{\min}/n$ to simplify further the last quantity and define:
$T_{1}=\sqrt{\frac{C_{3}}{2}}\frac{\mu^{2}\kappa^{2}L}{\|M\|_{\max}}\frac{r^{5/4}(m+n)^{3/4}(mn)^{1/4}}{\sqrt{m\rho_{\min}(m\wedge n)}}\sqrt{T}\log^{3/2}\left(1+mnT\right)\log\left(T(m+n)\frac{L^{2}}{\sigma^{2}}\right).$ (51)
This choice of $T_{1}$ entails that the regret can be upper bounded as follows:
$R^{\pi}(T)=\widetilde{O}\left(\mu^{2}\kappa^{2}r^{5/4}L\frac{(m+n)^{3/4}(mn)^{1/4}}{\sqrt{m\rho_{\min}(m\wedge n)}}\sqrt{T}\right).$
Step 4: (Checking conditions on $T_{1}$). It remains to check whether our choice of $T_{1}$ verifies $T_{1}\geq c\frac{L^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$ and $T_{1}\leq T$. These conditions are satisfied if
$T=\widetilde{\Omega}\left(\mu^{4}\kappa^{4}r^{5/2}L^{2}\frac{(m+n)^{3/2}\sqrt{mn}}{\|M\|_{\max}^{2}m\rho_{\min}(m\wedge n)}\right).$
Now, note that when the above condition does not hold, i.e.,
$T=\widetilde{O}\left(\mu^{4}\kappa^{4}r^{5/2}L^{2}\frac{(m+n)^{3/2}\sqrt{mn}}{\|M\|_{\max}^{2}m\rho_{\min}(m\wedge n)}\right),$
then the trivial bound $R^{\pi}(T)\leq 2T\|M\|_{\max}$, gives us after a few simple computations the upper bound
$R^{\pi}(T)\leq 2\|M\|_{\max}T=\widetilde{O}\left(\mu^{2}\kappa^{2}r^{5/4}L\frac{(m+n)^{3/4}(mn)^{1/4}}{\sqrt{m\rho_{\min}(m\wedge n)}}\sqrt{T}\right).$
If, furthermore, $M$ and $\rho$ are homogeneous, we obtain, for any $T\geq 1$:
$R^{\pi}(T)=\widetilde{O}\left(L(m+n)^{3/4}\sqrt{T}\right).$
Appendix H Reduction to almost low-dimensional linear bandits and its limitations
@@
We note that according to (5), $M_{i,j}=\langle\phi_{i,j}^{\textup{ext}},\theta^{\textup{ext}}\rangle$. Moreover, one can readily show that, for any $(i,j)\in[m]\times[n]$.
@@
$\|\theta_{\ell}\|_{2}\lesssim\frac{L^{2}\kappa r(m+n)}{T_{1}\omega_{\min}\sigma_{r}}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right)$
@@
From these observations, together with the reformulation $M_{i,j}=\langle\phi_{i,j}^{\textup{ext}},\theta^{\textup{ext}}\rangle$, one can claim a reduction from a low-rank bandit model to an almost low-dimensional linear bandit (Valko et al., 2014; Jun et al., 2019; Kocák et al., 2020; Kang et al., 2022). More precisely, the sub-vector corresponding to the last $(m-r)(n-r)$ components of $\theta$ (i.e., those corresponding to $\theta_{\ell}$) do not have a significant impact on the rewards, hence reducing the effective dimension of this linear bandit to $d=r(m+n)-r^{2}$.
We can derive an analogue of Theorem 8.1 with a similar regret analysis:
@@
$\Lambda=\mathrm{diag}(\overbrace{\lambda,\dots,\lambda}^{d},\overbrace{\lambda_{\perp},\dots,\lambda_{\perp}}^{mn-d}),\qquad\beta=\sigma\sqrt{2\log\left(\frac{10Tmn}{\delta}\log(1+T)\right)}+B_{2}\sqrt{\lambda}+B_{\ell}\sqrt{\lambda_{\perp}}.$
Then, Algorithm 2 with SupLinUCB parameters $T-T_{1},\beta(\delta),\Lambda$ has an expected regret that satisfies $R^{\pi}_{T}=\widetilde{O}\left(L\mu\kappa r\frac{\sqrt{m+n}(mn)^{1/4}}{\sqrt{m\rho_{\min}}}\sqrt{T}\right)$.
The proof of the proposition only differs slightly from the proof of Theorem 8.1. Instead of Theorem G.1, we combine the analysis of SupLinUCB [^19] Here we used SupLinUCB as given in (Chu et al., 2011) for the analysis. from (Chu et al., 2011), Lemma 38 in (Kocák et al., 2020), Lemma 3 (Jun et al., 2019), and Lemma 11 in (Abbasi-Yadkori et al., 2011) to show that the second phase of this algorithm has regret as follows:
${\cal R}^{\pi}(T)\lesssim\sqrt{dT}\left(r_{\max}+\sigma\sqrt{\log\left(eTK/\delta\right)}+B_{2}\sqrt{\lambda}+B_{\ell}\sqrt{\lambda_{\perp}}\right)\sqrt{\log\left(1+T/\lambda\right)\log T}$
@@
(i) First, it does not take advantage of the delocalization of the subspace recovery error, i.e., of the fact that this error spreads out along $m+n$ dimensions. This delocalization property is quantified through our subspace recovery error guarantees in the two-to-infinity norm, and it has subtle but important implications. Indeed, it can be leveraged to perform a reduction to an almost low-dimensional linear bandit that is also almost sparse, in the sense that $\|\theta_{\ell}\|_{\infty}\lesssim\|\theta_{\ell}\|_{2}/\sqrt{mn}$. A reduction to a misspecified linear bandit leverages this structure and allows us to devise algorithms with tighter regret guarantees.
@@
(iii) Finally, a reduction to misspecified linear bandit is more appealing computationally because the resulting feature vectors are of dimension $r(m+n)-r^{2}$. In the case of a reduction to an almost low-dimensional linear bandit, these vectors are of dimension $mn$. This means that instead of having to invert [^20] Indeed, almost all existing algorithms for linear and contextual linear bandits require inverting the features matrix during a regression step that involves least squares estimation (Lattimore & Szepesvári, 2020). $mn\times mn$ matrices, we will only have to invert $(r(m+n)-r^{2})\times(r(m+n)-r^{2})$ matrices.
@@ Appendix I Additional discussions
I.1 Knowledge of the context distribution
We decided to assume knowledge of $\rho$ to simplify the exposition of our work, as dealing with unknown $\rho$ does not add any insight to our work nor does it harm any of our guarantees, as we show in this section. After $T$ observations, a natural estimator $\hat{\rho}_{T}$ of $\rho$ is:
$\forall i\in[m],\qquad\hat{\rho}_{i,T}=\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\{i_{t}=1\}.$
Next, an immediate application of Bernstein’s inequality gives
$\mathbb{P}(|\hat{\rho}_{i,T}-\rho_{i}|>\varepsilon\rho_{i})\leq 2\exp\left(-\frac{T\rho_{i}^{2}\varepsilon^{2}}{2\sigma_{i}^{2}+2\varepsilon\rho_{i}/3}\right)\leq 2\exp\left(-\frac{T\rho_{i}^{2}\varepsilon^{2}}{2\rho_{i}+2\rho_{i}\varepsilon/3}\right)\leq 2\exp\left(-T\rho_{i}\left(\frac{\varepsilon^{2}}{2+2\varepsilon/3}\right)\right)$
where we used $\sigma_{i}^{2}:=\mathbb{E}[(\mathds{1}\{i_{t}=i\}-\rho_{i})^{2}]=\rho_{i}-\rho_{i}^{2}\leq\rho_{i}$. Using a union bound and choosing for instance $\varepsilon=1/2$, we obtain that
$\mathbb{P}(\exists i\in[m],\;|\hat{\rho}_{i,t}-\rho_{i}|>\rho_{i}/2)\leq 2\sum_{i=1}^{m}\exp\left(-\frac{3T\rho_{i}}{16}\right)\leq 2m\exp\left(-\frac{3T\rho_{\min}}{16}\right).$
Thus, for all $\delta\in(0,1)$,
$\mathbb{P}(\forall i\in[m],\;|\hat{\rho}_{i,t}-\rho_{i}|\leq\rho_{i}/2)\geq 1-\delta$
provided that
$T\geq\frac{16}{3\rho_{\min}}\log\left(\frac{2m}{\delta}\right).$
In view of the above result, estimating $\rho$ always requires fewer samples than required in Theorem 4.1. To see that, observe that by Lemma B.8 and the inequality $\omega_{\min}\leq\rho_{\min}/n$,
$\frac{L^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)\gtrsim\frac{m+n}{\omega_{\min}mn}\log\left(\frac{2m}{\delta}\right)\gtrsim\frac{1}{\rho_{\min}}\log\left(\frac{2m}{\delta}\right).$
Consequently, $\rho$ can be assumed to be known without loss of generality.
I.2 Low-rank matrix bandits
Our framework can be applied to address a related and simpler problem known as low-rank matrix bandits. In this scenario, the learner must choose two arms, $i_{t}\in[m]$ and $j_{t}\in[n]$, at each time step $t\in T$. In contrast, in the contextual low-rank bandit setting, the row $i_{t}$ is sampled according to a distribution $\rho$ that is not chosen by the learner.
@@ I.2 Low-rank matrix bandits
Similarly to Theorem 8.1, it can be proven that such an algorithm would achieve a regret of at most $\widetilde{O}(L\mu^{2}\kappa^{2}r^{5/4}(m+n)^{3/4}(mn)^{1/4}\sqrt{T/(m\wedge n)})$ under an appropriate condition on $T$ (see Theorem 8.1). When $M$ is homogeneous, the regret scales as $\widetilde{O}(L(m+n)^{3/4}\sqrt{T})$.
Most existing work on low-rank bandits lacks minimax regret guarantees. (Katariya et al., 2017) introduced Rank1Elim, achieving regret $\widetilde{O}((m+n)\log(T)/\tilde{\Delta}_{\min})$ [^21] Here $\tilde{\Delta}_{\min}$ corresponds to a notion of minimum gap different from the standard one - see (Katariya et al., 2017). for a rank-1 reward matrix. (Kveton et al., 2017) extended this result to rank $r$, with additional strong assumptions on the reward matrix in their LowRankElim algorithm. (Trinh et al., 2020) provided the first algorithm with asymptotically optimal instance-dependent regret for rank-$1$ bandits. (Bayati et al., 2022) proposed an algorithm enjoying a minimax regret guarantee, but their upper bounds scale at least with $mn$ through the constant $C_{2}$ in the worst case (see Theorem 2 in (Bayati et al., 2022)), and requires tuning a filtering resolution $h$, making their guarantees hard to evaluate. (Stojanovic et al., 2024) presented SME-AE algorithm, leveraging the entry-wise matrix estimation guarantees to obtain a regret upper bound of order $\widetilde{O}((m+n)(\bar{\Delta}/\Delta_{\min}^{2})\log^{3}(T))$ where $\bar{\Delta}$ is the average reward gap, and $\Delta_{\min}$ is the minimum reward gap.
While these gap-dependent bounds require strong assumptions, e.g., (Stojanovic et al., 2024) assume that $\Delta_{\max}/\Delta_{\min}\leq\zeta$ where $\Delta_{\max}$ is the maximum reward gap, we provide the first algorithm with a minimax regret of order $\widetilde{O}((m+n)^{3/4}\sqrt{T})$ without such assumptions.
I.3 Improved max-norm guarantees
@@
$\|M-\widehat{M}\|_{\max}\lesssim L\sqrt{\frac{\mu^{6}\kappa^{4}r^{3}(m+n)}{T\omega_{\min}(m\wedge n)^{2}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}$
with probability larger than $1-\delta$ when $T\gtrsim\frac{L^{2}\kappa^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$.
@@
with probability larger than $1-\delta$ under the assumptions of Proposition D.2, where $K$ depends polynomially on the model parameters. When $T$ is large enough to ensure that the first term in the error bound is larger than the second, the scaling in $\mu$ and $r$ of the max-norm error bound is improved from $\mu^{3}r^{3/2}$ to $\mu r^{1/2}$. Furthermore, the dependency in $\kappa$ and $\|M\|_{\max}$ is removed in the first term (recall that $L=\|M\|_{\max}\vee\sigma$). This may appear suprising, but note that the term $\widetilde{O}\left(\frac{K}{T}\right)$ still depends on $\kappa$ and $\|M\|_{\max}$ through $K$: more samples are required to achieve the stated scaling when $\kappa$ and $\|M\|_{\max}$ are large.
I.4 Removing the dependence in $\|M\|_{\max}$ in the regret upper bound
We assume that $m=\Theta(n)$ for simplicity as in (Chen et al., 2020). Let $\widehat{M}:=Z_{\mathrm{cvx},r}$ from Theorem 2 in (Chen et al., 2020) and let $\widehat{M}=\widehat{U}\widehat{\Sigma}\widehat{V}^{\top}$. Then, we have:
Lemma I.1.
Under the condition $\|M-\widehat{M}\|_{\textup{op}}\leq\sigma_{r}(M)/4$, it holds that:
$\max\left(\|U-\widehat{U}(\widehat{U}^{\top}U)\|_{2\to\infty},\|V-\widehat{V}(\widehat{V}^{\top}V)\|_{2\to\infty}\right)\lesssim\frac{1}{\sigma_{r}(M)}\left(\sqrt{n}\|M-\widehat{M}\|_{\max}+\mu\sqrt{\frac{r}{n}}\|M-\widehat{M}\|_{\textup{op}}\right)$
Proof.
Using repeatedly that $\|AB\|_{2\to\infty}\leq\|A\|_{2\to\infty}\|B\|_{\textup{op}}$ for appropriate $A,B$, we obtain:
$\|U-\widehat{U}\widehat{U}^{\top}U\|_{2\to\infty}$ $=\|MV\Sigma^{-1}-\widehat{U}\widehat{U}^{\top}U\|_{2\to\infty}\leq\frac{1}{\sigma_{r}(M)}\|MV-\widehat{U}\widehat{U}^{\top}U\Sigma\|_{2\to\infty}$
$\leq\frac{1}{\sigma_{r}(M)}\|MV-\widehat{M}V\|_{2\to\infty}+\frac{1}{\sigma_{r}(M)}\|\widehat{M}V-\widehat{U}\widehat{U}^{\top}U\Sigma\|_{2\to\infty}$
Note that the first term is bounded by $\sqrt{n}\|M-\widehat{M}\|_{\max}$ since $\|MV-\widehat{M}V\|_{2\to\infty}\leq\|M-\widehat{M}\|_{2\to\infty}\|V\|_{\textup{op}}\leq\sqrt{n}\|M-\widehat{M}\|_{\max}$. Before we bound the second term, note that:
$\widehat{U}\widehat{U}^{\top}U\Sigma=\widehat{U}\widehat{U}^{\top}MV=\widehat{U}\widehat{U}^{\top}(M-\widehat{M})V+\widehat{U}\widehat{U}^{\top}\widehat{M}V$
and $\widehat{U}\widehat{U}^{\top}\widehat{M}V=\widehat{U}\widehat{\Sigma}\widehat{V}^{\top}V=\widehat{M}V$. Thus:
$\|\widehat{M}V-\widehat{U}\widehat{U}^{\top}U\Sigma\|_{2\to\infty}=\|\widehat{U}\widehat{U}^{\top}(M-\widehat{M})V\|_{2\to\infty}\leq\|\widehat{U}\|_{2\to\infty}\|M-\widehat{M}\|_{\textup{op}}.$
Lastly, we bound $\|\widehat{U}\|_{2\to\infty}$ as follows:
$\|\widehat{U}\|_{2\to\infty}\leq\|\widehat{U}\widehat{U}^{\top}U\|_{2\to\infty}\|(\widehat{U}^{\top}U)^{-1}\|_{\textup{op}}\leq(\|U\|_{2\to\infty}+\|U-\widehat{U}\widehat{U}^{\top}U\|_{2\to\infty})\|(\widehat{U}^{\top}U)^{-1}\|_{\textup{op}}$
and similarly to the equation between (49) and (50) in (Stojanovic et al., 2024) we have:
$\|(\widehat{U}^{\top}U)^{-1}\|_{\textup{op}}=\frac{1}{\sigma_{r}(\widehat{U}^{\top}U)}\leq\frac{1}{1-\|\widehat{U}^{\top}U-\mathrm{sgn}(\widehat{U}^{\top}U)\|_{\textup{op}}}\leq\frac{1}{1-\frac{2\|M-\widehat{M}\|_{\textup{op}}^{2}}{\sigma_{r}^{2}(M)}}$
Under the condition $\|M-\widehat{M}\|_{\textup{op}}\leq\sigma_{r}(M)/4$ we consequently have:
$\|U-\widehat{U}\widehat{U}^{\top}U\|_{2\to\infty}\lesssim\frac{1}{\sigma_{r}(M)}\left(\|M-\widehat{M}\|_{2\to\infty}+\|U\|_{2\to\infty}\|M-\widehat{M}\|_{\textup{op}}\right)$
Then, as a consequence of Lemma I.1 and Theorem 2 from (Chen et al., 2020), we obtain analogously to Theorem 4.1:
Corollary I.2.
$\epsilon_{\textup{Sub-Rec}}=\widetilde{O}\left(\frac{n\sigma}{\sigma_{r}}\frac{1}{\sqrt{T}}\mu^{3}\kappa^{5/2}r^{3/2}\right)$
holds with probability $\geq 1-O(n^{-3})$, provided that $T\gtrsim\kappa^{4}\mu^{2}rn\log n\left(\mu^{2}r\log^{2}n+\frac{\sigma^{2}n^{2}}{\sigma_{r}^{2}}\right).$
Considering the setting of Corollary I.2, the bound from Theorem 4.1 scales with $\widetilde{O}\left(\frac{nL}{\sigma_{r}}\frac{1}{\sqrt{T}}\mu\kappa r^{1/2}\right)$. Note that even though the dependence on $L=\|M\|_{\max}\vee\sigma$ is reduced to $\sigma$ in Corollary I.2, the scaling in $\mu,\kappa$ and $r$ is worse than in Theorem 4.1.
To derive the corresponding regret bound, we can rewrite the requirement on the number of samples from Corollary I.2 as $T_{1}\gtrsim n\left(1+\frac{\sigma^{2}}{\|M\|_{\max}^{2}}\right)\mathrm{poly}(\kappa,\mu,r,\log n)$. Next, similarly to (51) in the proof of Theorem 8.1 we obtain that the regret is minimized for:
$T_{1}^{\star}\gtrsim\frac{\sigma}{\|M\|_{\max}}\sqrt{T}n^{3/4}\mathrm{poly}(\kappa,\mu,r,\log n)$
Combining this with the requirement on $T_{1}$, we set:
$T_{1}:=\max\left\{\frac{\sigma}{\|M\|_{\max}}\sqrt{T}n^{3/4},n\left(1+\frac{\sigma^{2}}{\|M\|_{\max}^{2}}\right)\right\}\mathrm{poly}(\kappa,\mu,r,\log n)$
Next, following the arguments from the proof of Theorem 8.1 in Appendix G.2 we obtain that combining our analysis with the guarantee from Corollary I.2 achieves:
$R^{\pi}(T)\lesssim\max\left\{\sigma\sqrt{T}n^{3/4},n\left(\|M\|_{\max}+\sigma\frac{\sigma}{\|M\|_{\max}}\right)\right\}\mathrm{poly}(\kappa,\mu,r,\log n),$
Comparing this result to the regret upper bound from Theorem 8.1, namely
$R^{\pi}(T)=\widetilde{O}\left((\|M\|_{\max}\vee\sigma)\mu^{2}\kappa^{2}r^{5/4}n^{3/4}\sqrt{T}\right),$
we see that in the regime $\sigma\lesssim\|M\|_{\max}$, the regret upper bound of Theorem 8.1 scales with $\|M\|_{\max}n^{3/4}\sqrt{T}$, whereas the regret upper bound obtained by using Corollary I.2 scales with $\sigma n^{3/4}\sqrt{T}+\|M\|_{\max}n$ up to $\mathrm{poly}(\kappa,\mu,r,\log n)$ terms.
[…]